# Optimistic Verifiable Training by Controlling Hardware Nondeterminism

### v2

@@ 5.2 Primitives
5.3 Trainer
The $\mathrm{trainer}$’s task begins when a client approaches them with dataset $D$, training specifications (epochs $E$, loss function $\mathsf{loss}$, etc.), and a requested model precision $b_{m}$.The $\mathrm{trainer}$ can then choose a training precision $b_{tr}>b_{m}$, a rounding amount $b_{r}\leq b_{m}$, and a checkpointing interval $k$ to periodically store small $\mathsf{hash}_{\mathsf{sha256}}(\theta)$ of model weights $\theta$ in a Merkle tree, for efficient comparison with an eventual $\mathrm{auditor}$. Then, as detailed in Algorithm 2, the $\mathrm{trainer}$ can perform training as normal, but after every intermediate computation (e.g. convolution) perform the $\mathsf{rnd}_{b_{r}}$ operation on each output. Rounding is applied to computations in both the forward and backward passes. Finally, either using a fixed threshold $\tau$ or a layer-specific optimal $\tau$ from the $\mathsf{threshold}$ function described in Section 5.5, the $\mathrm{trainer}$ applies $\mathsf{log}$, which logs rounding choices only for the computations an auditor should copy. The output of the algorithm includes a rounding log file $F$ and the root of the Merkle tree which, along with the shared randomness R and all training parameters, the $\mathrm{trainer}$ can share with any trusted third-pary $\mathrm{auditor}$ who challenges them.
5.4 Auditor
After a client approaches them (and shares their data), the $\mathrm{auditor}$ can initiate the verification game described in Section 3. To avoid facing a penalty, the $\mathrm{trainer}$ must cooperate by sharing the chosen rounding amount $b_{r}$, randomness R used in training (e.g. a pseudo-random number generator), details of the model architecture and training parameters, the checkpointing interval $k$, and set of rounding logs $F$. The $\mathrm{auditor}$ then follows the training procedure and corrects their rounding choice (e.g. up or down) to match those logged in $F$ using the $\mathsf{rev}$ operation, as detailed in Algorithm 3 in the Appendix. By correcting for each rounding choice mismatch with the $\mathrm{trainer}$ during the course of training, the $\mathrm{auditor}$ is able to prevent errors due to nondeterminism to accumulate. Therefore, the $\mathrm{auditor}$ can store the $\mathsf{hash}_{\mathsf{sha256}}(\theta)$ of model weights $\theta$ in a Merkle tree at interval $k$, knowing that if training was done correctly, the model weights should be identical to the $\mathrm{trainer}$’s at any timestep. The output of Algorithm 3 is the root of the $\mathrm{auditor}$’s Merkle tree, which they can use to compare with the $\mathrm{trainer}$’s root.
5.5 Reducing storage cost
Logging rounding decisions for every neural network layer output during training incurs a large baseline storage cost, and is the main limitation of our work. More formally, for dataset $D$, batch size $B$, number of training epochs $E$, and model layers $L_{\theta}$, the upper bound on the total storage cost for verifiable training with our method is:
[…]
where $o_{l,f}$ and $o_{l,f}$ represent the size of outputs of the forward pass and backward pass of layer $l$. Note that the log entries do not need to be kept around in the RAM and can be written straight to the disk. Moreover, this cost is a one-time cost incurred by the $\mathrm{trainer}$, who in our context is likely to be a powerful commercial provider with access to such storage capacity. Furthermore, as we later show in Section 6, for models with many linear layers like Transformer-based language models (e.g. GPT-2), where parameters significantly outnumber intermediate computations, this storage cost is significantly smaller than alternative approaches that require saving model weights (Jia et al., 2021). Nevertheless, we now describe our method for reducing storage cost by (i) efficiently encoding rounding logs and (ii) adaptive selection of the threshold $\tau$ to reduce the storage costs.
Efficient Encoding: Recall that $\mathsf{log}(x)$ only logs one of three possible values from the set ${0,1,2}$, as opposed to the floating point model weights. We pack sub-sequences of five log entries into a single byte via a fast GPU-based radix-3 to radix-2 conversion, yielding 1.6 bits/entry storage that is close to the best possible packing of 1.58 bits/entry, and yields a 77% storage reduction in practice relative to naively storing one log entry per byte.
Adaptive Threshold: Recall that we need to select a threshold $\tau$ that controls for whether the $\mathrm{trainer}$ logs a rounding choice, or instead logs 1 which the $\mathrm{auditor}$ ignores. The more one can increase $\tau$, the more 1 values are recorded which, if significant, can make the rounding more compressible (due to an increase in long sequences of 1s). Furthermore, it is possible that the divergence $d_{div}$ between outputs on two different GPUs, given the same input, is function-specific. For example, while the convolution operation requires several matrix multiplications that might result in large floating point accumulation error, operations such as scaling are unlikely to result in large $d_{div}$, and a larger $\tau$ can be applied. We develop an efficient algorithm (Algorithm 4 in the Appendix) to find the optimal value for $\tau$ given a particular layer and data of output values that led to different rounding choices between any two GPUs (e.g. Case A in Figure 4). For a given rounding amount $b_{r}$ and training precision $b_{tr}$, the algorithm performs a binary search between $\tau=0.25\cdot\epsilon_{32}$ (our assumed upper bound for the $d_{div}$ between two GPUs for any function) and $\tau=0.5\cdot\epsilon_{b_{r}}$ (the rounding boundary). By performing this procedure for the different intermediate computations in a model, the $\mathrm{trainer}$ can hope to log more values of 1, leading to a stronger compression of the rounding log $F$.
Merkle Tree Storage: The only other additional storage required of our method is the Merkle tree of model checkpoints (which is negligible compared to the rounding logs). Storing SHA-256 hashes of model weights during training in a Merkle tree creates an efficient mechanism for the verification game described in Section 3, and the audit ends when either the $\mathrm{trainer}$ withdraws, the $\mathrm{auditor}$ confirms that training was performed correctly, or the $\mathrm{auditor}$ can present paths to the two leaves of their Merkle tree where the training runs first started to diverge, thereby providing evidence to dispute the $\mathrm{trainer}$. Comparing two Merkle trees given their root hash is a standard binary search procedure.
6 Empirical Results
ResNet-50 $b=32$ ResNet-50 $b=26$ GPT-2 $b=32$ GPT-2 $b=26$
Naive Encoding 456 MB 456 MB 92 MB 92 MB
Efficient Encoding 105 MB 105 MB 22 MB 22 MB
+ Zip Compression 96 MB 91 MB 20 MB 18 MB
Table 3: Average number of $\mathsf{rev}$ corrections performed by $\mathrm{auditor}$ per step of training. Even at $b=32$, the auditor only makes 20-25 corrections (2e-6 to 9e-6% of samples) per training step, suggesting significant opportunity for reducing the storage cost of our method.
ResNet-50 $b=32$ $b=31$ $b=30$ $b=29$ $b=28$ $b=27$ $b=26$
Forward $15\pm 3$ $6\pm 2$ $3\pm 1$ $3\pm 1$ 0 0 0
Backward $10\pm 0.6$ $6\pm 0.6$ $2\pm 1$ $0.7\pm 0.7$ $0\pm 0$ $0\pm 0$ $0\pm 0$
GPT-2 $b=32$ $b=31$ $b=30$ $b=29$ $b=28$ $b=27$ $b=26$
Table 4: Adaptive thresholds identified for different operations using the binary search procedure from Algorithm 4 with $b=32$.
2D Convolution Batch Norm Linear Layer Norm
Dimension 256 (1,1) (128, 128, 16, 16) (768,768) (768,1)
$\tau$ $0.305*2^{-23}$ $0.499*2^{-23}$ $0.465*2^{-23}$ $0.499*2^{-23}$
We evaluate our verifiable training method on the two large-scale models listed below with all possible $\mathrm{trainer}$ and $\mathrm{auditor}$ pairs across the following three NVIDIA GPU architectures: A40, TITAN Xp, and RTX 2080 Ti (see Appendix A for more details). Additionally, in Section 6.4, we compare our method with a proof-based system for logistic regression as no known proof-based systems exist for the large-scale models we have considered.
  1. 1.
    ResNet-50: We train (from random initialization) ResNet-50 (23M) on CIFAR-10 with dataset size 50,000 and batch size B=64. Training for 100 epochs results in test accuracy of 90.7% on a Titan RTX Ti.
  2. 2.
    GPT-2: We finetune GPT-2 (117M) on a corpus of Shakespeare text with dataset size 1.1M tokens, batch size B=8, and sequence length 64. Training for 1 epoch results in perplexity 4.22 on a Titan RTX Ti.
In Figure 2 we show that nondeterminism due to GPU architecture exists for both training tasks. While using library flags and deterministic settings allow us to repeatedly obtain identical results across training runs on the same GPU architecture, training on different GPU architectures results in fundamentally different models.
6.1 Implementation details
We implement our verifiable training method entirely on top of the pytorch framework, with torch version 1.13.1 and CUDA version 11.7. The intermediate computations we apply $\mathsf{rnd_{b}}$ to are layers (e.g. torch.nn.Conv2D) in the model’s computation graph. Rounding-related operations ($\mathsf{rnd}$ and $\mathsf{rev}$) either using casting or floating-point functions (e.g. torch.nextafter) that can run on the GPU, thus having little impact on computational speed. Because we observed that the torch.randn operation used for dropout in GPT-2 is non-deterministic for long inputs (even for the same seed, see Appendix E), we implement our own dropout as our method requires shared randomness $R$.
6.2 Successful control for non-determinism
Our method completely eliminates non-determinism between full training runs of both for both the ResNet-50 training and GPT-2 fine-tuning tasks across all possible $\mathrm{trainer}$ and $\mathrm{auditor}$ pairs between the A40, Titan XP, and RTX 2080 Ti GPUs. As Figure 5 shows, standard FP32 training results in an increasing divergence (l2-distance of weights) between models on different GPUs over the course of training. Furthermore, we show the simple approach of training in FP64 and rounding to FP32 after every intermediate computation, but without the $\mathrm{auditor}$ correcting rounding decisions with $\mathsf{rev}$, fails to mitigate this issue. Only our method, in which the $\mathrm{auditor}$ follows the rounding decisions ($b_{r}=32$) made by the $\mathrm{trainer}$ for every intermediate computation, eliminates non-determinism and persists over time.
Our implementation, which requires disk I/O during training to store the rounding decisions, results in a small increase in training time for the $\mathrm{trainer}$ (1.2-1.4x) and $\mathrm{auditor}$ (1.3-1.7x) using a non-optimized, protoype implementation (Table 5). We report the storage requirements of our method in Table 2, showing that our efficient encoding scheme reduces the size of the $\mathrm{trainer}$’s rounding logs by 77%, relative to naive logging. Because the Merkle tree stores 32-byte SHA-256 hashes, its overall size (KBs) and creation time are negligible and not reported.
Finally, we show that decreasing the rounding amount $b$ to values even as low as 26 has little effect on model performance (we observe no change in accuracy, so report test loss), but increase training time (Figure 5). We observe that smaller values of $b$ do allow more log entries to be ignored, improving compression of the file, which we discuss next.
6.3 Compression with adaptive threshold
Our approach particularly outperforms (Table 2) the storage costs of proof-of-learning protocols that save model weights for GPT-2 (2.78GB), which has many linear layers – we observe more than 140x reduction relative to the approach in Jia et al. (2021). We further reduce the storage cost of our method by decreasing the rounding amount $b$ and implementing the adaptive thresholding strategy (Section 5.5).
In Table 4, we report adaptive thresholds $\tau$ for 4 different pytorch layers at rounding amount $b_{r}=32$. Convolutions require the lowest $\tau$, indicating larger divergence in outputs between GPU types, which is expected due to the large number of matrix multiplications. Meanwhile, $\tau$ is higher for normalization layers, likely due to smaller divergences between GPU types. Because adaptive thresholding seeks to reduce the number of times rounding decisions (0 and 2) are logged and improve log file compression, we report storage cost after zip compression in Table 2. As expected, more aggressive rounding (which results in a higher $\tau$) improves the compression rate. Although the compression gains are mild in comparison to our encoding step, such differences build-up over the course of training.
Finally, we report the average number of $\mathsf{rev}$ corrections an $\mathrm{auditor}$ needs to perform for one training step in our two tasks (Table 3). These values are surprisingly small in comparison to the number of operations logged – only a maximum of 25, or 2e-6% (ResNet-50) and 9e-6% (GPT-2) of logged values, are actually needed by the $\mathrm{auditor}$! We also observe that severe rounding (e.g. $b=27$) completely eliminated the hardware non-determinism for our tasks, requiring no corrections from the $\mathrm{auditor}$. This shows a huge gap between the # of values currently saved by the $\mathrm{trainer}$ and those actually needed by the $\mathrm{auditor}$. An exciting direction for future work, therefore, is to develop methods that could reliably predict when a divergence will not occur and skip logging entirely, reducing the overall storage cost of our method.
6.4 Comparison with alternative approaches
Logistic Regression Garg et al. (2023) recently proposed a zero-knowledge proof-based system for verifiable training of a logistic regression, which importantly does not leak information about the client’s data or require a trusted third-party $\mathrm{auditor}$, unlike our work. However, since verifiable training itself is motivated by a client not having sufficient resources to train the model, it is crucial to consider the implications of scale. The authors report the prover time and proof size requirements for one training pass of logistic regression on a dataset of $2^{18}$ items, with 1024 dimensions and a batch size of 2014, as 72 seconds (training and proof generation time) and 350 MB respectively. We replicate this training task, and find that our method significantly improves upon both storage and time requirements, requiring only 106 KB and 7 seconds (both training and auditing). Furthermore, because Garg et al. (2023) do not report the duration of “offline phase” of their method, their reported value is a lower bound on the actual time required. Finally, we note that the original proof-of-learning protocol from Jia et al. (2021), which also considers a trusted third-party, would require 9.2 MB per training step to store all model weights, and therefore our method is the most space efficient by a factor of at least 85x.
VGG-11 Concurrent to our work, Abbaszadeh et al. (2024) introduce a zero-knowledge proof-of-training protocol for deep neural networks, presenting results for one batch step of training for a simplified version of the VGG-11 model with 10M parameters, which is less than the original VGG-11 network and ResNet-50 (Simonyan & Zisserman, 2015). While the authors unfortunately do not provide the exact details of the simplified architecture, we can assume that increasing the number of parameters to the original VGG-11 would only increase their reported proof time and size. We therefore compare their reported values with an implementation of our method for the same task of verifying training of VGG-11 on CIFAR-10 with a batch size of 16. While their use of incrementally verifiable computation leads to tractable proof size (1.36MB vs. the 1.2MB per iteration cost of our method), Abbaszadeh et al. (2024)’s method requires 22 minutes per training iteration. In comparision, our method requires training and auditing times of only 6 seconds per iteration, and is thus more efficient by a significant factor of 220x, which is an important consideration when model training is provided as a commerical service.
Secure GPT-2 Inference The previously discussed proof-based systems for verifiable training by-pass the need for a third-party $\mathrm{auditor}$, but very few efficient systems exist in the literature. Many more works study secure inference of deep neural networks, which could be used to construct verifiable training protocols with stronger security guarantees than ours (e.g., allowing a $\mathrm{trainer}$ to keep a proprietary model’s weights private), but come at a significant cost to performance and resources. To demonstrate this, we consider adapting Gupta et al. (2023)’s protocol for secure inference of GPT-2 based on multi-party computation, to our context of verifiable training. Gupta et al. (2023) show how two parties, the $\mathrm{client}$ with private data and the $\mathrm{trainer}$, can jointly compute the forward pass of a known model architecture without revealing additional information beyond the model output to each other. Because they report the the communication overhead $P=0.37$GB and time $T=0.96$ seconds for one forward pass on a single data input, we can calculate $2\times P\times D\times E=\textbf{ 189 GB}$ and $2\times T\times D\times E=$ 983 seconds as estimated communication cost and time, respectively, for 1 step of training in out GPT-2 task, where 2 considers both the forward and backward pass. Compared with our method’s required storage cost (18MB) and training time (11s for training, 13.5 seconds for auditing), scaling Gupta et al. (2023)’s protocol for training would introduce around a 10,000x data and 40x time overhead.
We acknowledge that the previous comparisons are not straightforward, as they focus on providing strong security guarantees that do not require a trust third-party. However, this comes at the cost of relying on cryptographic techniques that are difficult to scale, and approximating non-linear functions (e.g. softmax in Transformer attention). By highlighting the trade-offs such methods provide, we hope to motivate future verifiable training schemes that, if one is willing trust a third-party $\mathrm{auditor}$, are more practical for real-world training tasks.
7 Limitations and Future Work
Our scheme for verifiable training successfully controls for hardware nondeterminism and expands the pool of potential auditors of a model training service, allowing us to envision a world where a $\mathrm{client}$ can even use two competing service providers it trusts to audit each other. However, one main limitation of our work in comparison with proof-based systems is the need for all parties to trust this third-party $\mathrm{auditor}$. If the $\mathrm{trainer}$ is providing finetuning services on top of models that are not open-source (e.g. OpenAI), then our scheme will only work for the third-party auditors that the $\mathrm{trainer}$is willing to share their model weights.
Another limitation of our work is the storage cost. While we have shown that it is significantly smaller than alternative methods, the vast majority of stored rounding decisions are not used by the $\mathrm{auditor}$ and are therefore unnecessary (Section 6). Therefore, an exciting direction for future work is to figure out how to mitigate this gap by better predicting when divergence between GPU computations occurs. Recent work has similarly argued that a stronger profile of noise during the training process is necessary for robust verifiable training (Fang et al., 2023). Finally, other directions for future work also include adapting our method for distributed training, as well as considering adversarial manipulation of the rounding logs by the trainer.
8 Acknowledgements
We would like to thank Bill Dally and Duncan Riach for a helpful discussion at the early stage of the project. Megha Srivastava was supported by an IBM PhD Fellowship and the NSF Graduate Research Fellowship Program under Grant No. DGE-1656518. In addition, this work was funded by NSF, DARPA, the Simons Foundation, UBRI, and NTT Research. Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA.
Appendix A GPU Details
  • NVIDIA Titan XP: 3840 Cores, 12 GB
  • NVIDIA RTX 2080 Ti: 4352 Cores, 11 GB
  • NVIDIA A40: 10752 Cores, 48 GB
Appendix B Audit Algorithm
See Algorithm 3.
Appendix C Adaptive Thresholding Algorithm
See Algorithm 4.
Appendix D Time Requirements
See Table 5.
Appendix E Random Number Generation
Our verifiable training scheme requires shared randomness between the $\mathrm{trainer}$ and $\mathrm{auditor}$, which is used for deciding input data batching, weight initialization, and operations such as dropout (randomly setting outputs to zero). More formally, our scheme requires sharing the same random seed and pseudo-random generator. However, in our implementation based on pytorch (assuming the same software version between $\mathrm{trainer}$ and $\mathrm{auditor}$), we chose to rely on the the torch random seed functionality. While this successfully controls for batch input ordering and weight initialization, it is unfortunately not sufficient for random number generation, as operations such as torch.nn.randn() leverage parallelism when the requested number of values is higher than a certain amount. Specifically, we found that across T40, RTX 2080 Ti, V100, A40, and A100, given the same seed, torch.randint() produces identical tensors onlt up to size 40960. At size 40961, T40 (which is an older GPU) deviated from the rest. Likewise, at size 69633, 2080 Ti deviated from the rest, and so on. Based on these observations, we arranged for calls to torch.randint() in the dropout layer (which is the only operation using large random tensors in our tasks) to be replaced by generating and concatenating multiple random tensors of size 40960 or less. Specifically, a random tensor of size $n>40960$ is generated by concatenating ($n//40960$) random tensors of size 40960 and one random tensor of size ($n\%40960$). However, we emphasize that it is therefore important in our scheme either for both parties to implement this change a priori, or simply use an external source for pseudorandomness.
@@ Appendix E Random Number Generation
2:  $N,T\leftarrow$ initialize large number of data points and iterations
@@
5:     $x\leftarrow$ select random input for layer $l$ in $b_{tr}$ floating point precision

### v3

@@ Appendix I Random Number Generation
2:  $N,T\leftarrow$ initialize large # of data points and iterations
@@
5:     $x\leftarrow$ select random input for layer $l$ in $b_{tr}$ floating-point precision
@@
Table 6: Comparison of model divergence due to data ordering versus GPU non-determinism. Reported numbers are averaged between 10 pairs of models, error bars are standard deviation.
Metric Data Ordering GPU Non-determinism
l2 weight difference $133.2\pm 9$ $1.1\pm 0.07$
l2 output distance $5.3\pm 0.03$ $0.26\pm 0.02$
Algorithm 4 $\mathsf{log}$
Input: value $x$, rounding amount $b_{r}$, threshold $\tau$, file $F$
1:  $\exp\leftarrow$ get exponent of $x$
2:  if $|x-\mathsf{rnd}_{b_{r}}(x)|>\exp*\tau$ $x<\mathsf{rnd}_{b_{r}}(x)$ then
3:     $\mathsf{write}(2,F)$ // log rounding up
4:  else if $|x-\mathsf{rnd}_{b_{r}}(x)|>\exp*\tau$ $x>\mathsf{rnd}_{b_{r}}(x)$ then
5:     $\mathsf{write}(0,F)$ // log rounding down
6:  else
7:     $\mathsf{write}(1,F)$ // log rounding ignore
8:  end if
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer:[Yes]
  9. Justification: See Limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer:[Yes]
  19. Justification: See Experiments and Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code will be released in public version.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Experiments and Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: No experiment results have variation necessitating statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix for GPUs.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Security Analysis.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: [N/A]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.